OpenStudy (anonymous):

What is the value of the x variable in the solution to the following system of equations? 4x − 3y = 3 5x − 4y = 3 x can be any number as there are infinitely many solutions to this system There is no x value as there is no solution to this system −3 3

OpenStudy (merchandize):

x=3, and y=3. Given your choices, I simply substituted in. 12-3y=3 -3y=-9 y=3 and 15-4y=3 -4y=-12 y = 3